OpenStudy (anonymous):

A car passes a landmark on a highway traveling at a constant rate of 55 kilometers per hour. One hour later, a second car passes the same landmark traveling in the same direction at 63 kilometers per hour. How much time after the second car passes the landmark will it overtake the first car? Do not do any rounding.

OpenStudy (anonymous):

Okay. First, think about what conditions mean that one car passes another...

OpenStudy (anonymous):

I don't understand?

OpenStudy (anonymous):

well their distance travelled after the landmark must be same right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay so then consider that the distance is x, alright? Then you know their speeds and the time difference. The second car starts one hour later. So, whatever time in which they reach at that same point, value of t for second car is t-1 hrs. Now equate speed/time for both

OpenStudy (anonymous):

I don't know how to do that. D:

OpenStudy (anonymous):

I'm sorry, I really don't understand this. I need help.

OpenStudy (anonymous):

Refer to the attachment from Mathematica.

OpenStudy (anonymous):



OpenStudy (anonymous):

following?

OpenStudy (anonymous):

Yes sir/ma'am

OpenStudy (anonymous):

Then, distance=speed*time, right? That would mean 55*t=63*(t-1)

OpenStudy (anonymous):

solve this for t!

OpenStudy (anonymous):

7.875?

OpenStudy (anonymous):

yeah, sorry but the question need t-1 right?

OpenStudy (anonymous):

I chose t to be time elapsed for first... so 6.875 will be the answer.

OpenStudy (anonymous):

Oh, I understand now! Thanks. I have 5 more questions like this to do and I think I can get them on my own now. :)

OpenStudy (jhannybean):

shouldn't it be `t+1` since it was an hour later?

OpenStudy (jhannybean):

`t-1` would mean the second car left an hour before the first, no?

OpenStudy (anonymous):

I chose the other value for t for the question @Jhannybean. and no an hour late would mean lesser time elapsed right?.

OpenStudy (anonymous):

t represents time elapsed after crossing the landmark.

OpenStudy (jhannybean):

time it takes for the first car to pass the landmark = t time it takes for the second car to pass the landmark an hour later = t+1

OpenStudy (anonymous):

Let t be the time to catch the first car. 55 (1 + t) = 63 t // Solve t = 55/8 = 6.875

OpenStudy (anonymous):

Thank you for the medal.